The perimeter of a rectangle is 120 feet. Use an equation to find the length and width of the rectangle if the length is twice the width.
Use P = 2W + 2L.

Answer by mananth(16946)

let width be x
width will be 2x
..
perimeter = 2L +2W
2*2x+2*x=120
6x+2x=120
/6
x= 20 feet the width
length = 2*20 = 40 feet
...
CHECK
20+20+40+40=120
...
m.ananth@hotmail.ca